A business firm sells goods at the price of Rs 450.The firm has decided to reduce the price of good to Rs 350.Consequently, the quantity demanded for the good rose from 25,000 units to 35,000 units. Calculate the price elasticity of demand.
1
Expert's answer
2020-11-30T16:37:43-0500

P1 = 450

P2 = 350

Q1 = 25000

Q2 = 35000

Percentage change in quantity "= \frac{Q_2 – Q_1}{Q_1} \times 100"


"= \frac{35000 – 25000}{25000} \times 100"


"= \frac{10000}{25000} \times 100 = 40" %


Percentage change in price "= \frac{P_2 – P_1}{P_1}\times 100"


"= \frac{350 – 450}{450}\times 100"


"= \frac{-100}{450}\times 100 = -22.22" %


Price elasticity of demand "= \frac{Percentage \;change\; in \;quantity}{Percentage \;change \;in\; price}"


"= \frac{40}{-22.22} = -1.8"


The price elasticity of demand is -1.8. The negative sign shows an inverse relationship between price and quantity demanded.